the sum of the lengths of the two legs of a right angle triangle is 14 cm. find the length of the shortest hypotenuse.
Answer by checkley71(8403)

THE SHORTEST HYPOTENUSE WOULD BE WITH AN ISOSCELES TRIANGLE (2 EQUAL SIDES OF 7 CM EACH).
7^2+7^2=C^2
49+49=C^2
98=C^2
C=SQRT98
C=9.8995 CM FOR THE HYPOTENUSE.
PROOF:
7.1^2+6.9^2=C^2
50.41+47.61=C^2
C^2=98.02
C=SQRT98.02
C=9.9 CM FOR THIS HYPOTENUSE.